OpenStudy (freeg13):

Jenna buys 12 lb of meat. Some of the meat is chicken and costs $1.90/lb. The rest is beef, which costs $4.30/lb. She spends a total of $34.80. Which equation models this situation? Let c represent the number of pounds of chicken she buys. A. (1.9 + 4.3)(12 – c) = 34.8 B. 4.3c + 1.9(c + 14) = 34.8 C. 4.3c + 1.9(12) = 34.8 D. 1.9c + 4.3(12 − c) = 34.8

OpenStudy (anonymous):

c is the pounds of chicken at $1.90/lb. How can you write the total cost of the chicken?

OpenStudy (freeg13):

well C is the amount of chicken there is

OpenStudy (freeg13):

or the pounds of chicken she buys

OpenStudy (deadpoolroxz):

@Vocaloid

OpenStudy (anonymous):

c is the pounds of chicken she buys. And she pays $1.90 for each pound. So the total amount she pays for chicken is the 1.90 * c

OpenStudy (freeg13):

but your forget beef

OpenStudy (anonymous):

didn't forget. Just making sure you got that part first

OpenStudy (anonymous):

So if 12 pounds of meat total was bought, and c pounds of it was chicken, how can you write the amount of beef sold?

OpenStudy (anonymous):

beef = all meat - chicken

OpenStudy (nikato):

well the total of chicken and beef is 12 pounds let pounds of chicken=c and let pounds of beef=b so c+b=12 ok?

OpenStudy (freeg13):

ok

OpenStudy (anonymous):

right, so then solve that for b

OpenStudy (nikato):

and you know it cost 1.9 per pound of chicken, which you can write as 1.9c and 4.3 per pound of beef, which is the same as 4.3b and you know the total cost is 34.80. can you try to make an equation using these three information?

OpenStudy (freeg13):

C?

OpenStudy (yolomcswagginsggg):

I was called?

OpenStudy (anonymous):

how did you get C? what did you do?

OpenStudy (freeg13):

i guessed

OpenStudy (nikato):

not exactly given that info. you're equation would be 1.9c + 4.3b=34.80 but remember our earlier equation. c+b=12 we want to isolate c so we have c=12-b now substitute this into the other equation

OpenStudy (yolomcswagginsggg):

D 100%

OpenStudy (freeg13):

oh thanks

OpenStudy (yolomcswagginsggg):

Np

OpenStudy (nikato):

yup^ i agree